In a three-digit number, the digit at the hundreds place is three times the digit at the ones place and the sum of the digits is 15. If the digits are reversed, the number is reduced by 396. Find the number (Ans. 672)
Answer by lwsshak3(11628)

In a three-digit number, the digit at the hundreds place is three times the digit at the ones place and the sum of the digits is 15. If the digits are reversed, the number is reduced by 396. Find the number (Ans. 672)
**
let u=units digit
let t=tens digit
let h=hundreds digit
..
h=3u
u+t+h=15
original number: 100h+10t+u
reversed number:100u+10t+h
original number-reversed number=396
(100h+10t+u)-(100u+10t+h)=396
100h+10t+u-100u-10t-h=396
99h-99u=396
99(3u)-99u=396
2u*99=396
u=396/2*99=2
h=3u=6
t=15-t+h=15-8=7
ans:
number: 672